OpenStudy (anonymous):

Find an equation of the line that is tangent to the graph of f and parallel to the given line. Function Line f(x) = 2x^2 2x − y + 5 = 0

OpenStudy (irishboy123):

the point is not to find the intersection of the curve and line rather it is to find a **new line** which: (A) is a tangent line to f(x); and (B) runs parallel to 2x − y + 5 = 0 this means solving \(f'(x) = m\) where m is the slope of the line 2x − y + 5 = 0. this should lead to a point \( (a, 2a^2)\) on the new line. then finish off by using point \((a,2a^2)\) & slope m to solve \(y = mx + c\) for the new line